OpenStudy (anonymous):

the perimeter of a rectangle is 32 ft. the length is 4 ft. longer than its width. find the length and width

OpenStudy (anonymous):

I have with 16 and 8 so far

OpenStudy (anonymous):

You can answer this question by setting up a system of equations.

OpenStudy (anonymous):

2x+2y then what

OpenStudy (anonymous):

2x+2y=32

OpenStudy (anonymous):

x=length y=width

OpenStudy (anonymous):

Lets call x your length and y your width. You know that x=y+4

OpenStudy (anonymous):

16+4?

OpenStudy (anonymous):

So far you have: 2x+2y=32 x=y+4

OpenStudy (anonymous):

Use a method known as substitution. Plug in x from your second equation into your first equation

OpenStudy (anonymous):

Do you understand how to do that?

OpenStudy (anonymous):

not yet looking at it 32+4?

OpenStudy (anonymous):

Try this: 2(y+4)+2y=32 2y+8+2y=32 4y=24 y=6

OpenStudy (anonymous):

So what I did there was I substituted x=y+4 into my first equation, 2x+2y=32

OpenStudy (anonymous):

I see what you did but what is 4y=24 and y=6?

OpenStudy (anonymous):

I combined 2y+2y (4y) and subtracted 8 from 32 (24)

OpenStudy (anonymous):

Once you have 4y=24, you can divide both sides by 4 in order to solve for y

OpenStudy (anonymous):

y=6?

OpenStudy (anonymous):

Yeah. And now you can use y to solve for x. Remember that x=y+4

OpenStudy (anonymous):

Length = 10 width = 6 would this be correct then?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

I have to go over this again from the top to double check our answers. Thank you very much. I want to make sure I understand.

OpenStudy (anonymous):

I checke "best answer", hope you received a medal

OpenStudy (anonymous):

No problem, if you have any questions just ask

OpenStudy (anonymous):

ok